OpenStudy (rebeccapink6398):

Kirk's family drove 36 miles in 50 minutes, which was 48% of the distance to his grandmother's house. What was the total distance to Kirk's grandmother's house? A. 48 miles B. 24 miles C. 17.28 miles D. 75 miles could you please help again @hoblos

OpenStudy (anonymous):

since 48% of total distance x= 36 miles, you can say /\[\frac{ 48 }{ 100 } * X= 36\]

OpenStudy (anonymous):

Now we only need to solve for X. so we can cross multiply, 48*X=36*100 48X=3600 x=3600/48 X=?

OpenStudy (rebeccapink6398):

75!?

OpenStudy (anonymous):

Yes! :)